# Grasping a Handful: Sequential Multi-Object Dexterous Grasp Generation

### v2

@@ IX Conclusion
We proposed SeqGrasp, an algorithm for sequentially grasping multiple objects with a dexterous hand. SeqGrasp combines OSes and differentiable-force closure to generate stable grasps that maximize the hand’s remaining DoF after each grasp. Using SeqGrasp, we constructed SeqDataset, currently the largest sequential grasping dataset, comprising 870K validated grasps across 509 diverse objects. This dataset enabled the training of SeqDiffuser, our diffusion-based sequential multi-object grasp sampler. The experimental evaluations demonstrated that SeqGrasp and SeqDiffuser outperformed the simultaneous multi-object grasping baseline MultiGrasp, achieving an 8.71%-43.33% higher average success rate. Moreover, SeqDiffuser proved to be 750-1250 times faster at generating grasps than SeqGrasp and MultiGrasp. In conclusion, this work demonstrated a fast and stable sequential multi-object grasp generation solution, which we hope can pave the way for more research in multi-object grasping.

### v3

@@ IX Conclusion
We proposed SeqGrasp, an algorithm for sequentially grasping multiple objects with a dexterous hand. SeqGrasp combines OSes and differentiable-force closure to generate stable grasps that maximize the hand’s remaining DoF after each grasp. Using SeqGrasp, we constructed SeqDataset, currently the largest sequential grasping dataset, comprising 870K validated grasps across 509 diverse objects. This dataset enabled the training of SeqDiffuser, our diffusion-based sequential multi-object grasp sampler. The experimental evaluations demonstrated that SeqGrasp and SeqDiffuser outperformed the simultaneous multi-object grasping baseline MultiGrasp, achieving an 8.71%-43.33% higher average success rate. Moreover, SeqDiffuser proved to be 750-1250 times faster at generating grasps than SeqGrasp and MultiGrasp. In conclusion, this work demonstrates a fast and stable solution for generating sequential multi-object grasps, which we hope will pave the way for further research in multi-object grasping.